water must be added t 20 millileters of a 4% minoxidle solution to dilute it to a 2% solution.how many milliters of water should be used
Answer by htmentor(1343)

The initial volume of minoxidil = (0.04)(20 ml) = 0.8 ml
Let x = the number of ml of water needed
A 2% solution will have 0.8 ml of minoxidil and x + 20 ml of solution:
0.8/(x+20) = 0.02
x+20 = 0.8/0.02
Solving for x gives x = 40 - 20 = 20
So 20 ml of water should be used